In a past​ election, the voter turnout was 69​%. In a​ survey, 934 subjects were asked if they voted in the election. a. Find the mean and standard deviation for the numbers of voters in groups of 934. b. In the survey of 934 ​people, 695 said that they voted in the election. Is this result consistent with the​ turnout, or is this result unlikely to occur with a turnout of 69​%? Why or why not? c. Based on these​ results, does it appear that accurate voting results can be obtained by asking voters how they​ acted?

a. μ = _______ (Round to one decimal place as​ needed.)

σ =_______ (Round to one decimal place as​ needed.)

b. Is the result of 695 voting in the election usual or​ unusual? (Choose one)

A). This result is usual because 695 is within the range of usual values.

B). This result is unusual because 695 is greater than the maximum usual value.

C). This result is unusual because 695 is within the range of usual values.

D). This result is unusual because 695 is below the minimum usual value.

c. Does it appear that accurate voting results can be obtained by asking voters how they​ acted? (Choose one)

A). No, because it appears that substantially more people say that they voted than the proportion of people who actually did vote.

B). No, because it appears that substantially fewer people say that they voted than the proportion of people who actually did vote.

C). Yes, because the results indicate that 69 % is a possible turnout.
Answer #1

Here we can use binomial distribution:

mean = mu = n*p = 934*0.69 = 644.5

Standard deviation = σ = Vn * p * (1-p) = V934 * 0.69 * 0.31 = 14.1

b) Here 695 is greater than the mean value 644.5

So we need to find maximum usual range = μ + 2 * σ 644.5 + (2 * 14.1) = 672.7

Since 695 is greater than 672.7 so correct option is "B). This result is unusual because 695 is greater than the maximum usual value."

c. Does it appear that accurate voting results can be obtained by asking voters how they​ acted?

From part b) the correct choice of part c) is  

"A). No, because it appears that substantially more people say that they voted than the proportion of people who actually did vote."